# Neural-Enhanced Born Approximation for Inverse Scattering

### v1

@@ 5.1 In-distribution performance
Following the discussion in Section 3.1, we generated 4,000 in-distribution scatterers to evaluate model performance. Figure 5 shows the actual profile of one example of a test scatterer together with the reconstructions.
To make clear that the CNN based models outperform the inverse Born approximation by itself, Figure 6 visualizes the error distribution for the test samples. The results show that the three models–CNN, CNNB, and BCNN–significantly outperform the Born approximation in terms of accuracy, reducing the relative $L^{2}$ error by almost 50%.
@@
Figure 5: An example of the reconstructions of three circular scatterers determined by each model. The ground truth is shown in the upper left panel. All CNN based methods are much more accurate than the inverse Born approximation since in this example the scattering is strong. CNNB (bottom left) shows the least speckle in the background.
Figure 6: The distribution of relative $L^{1}$ and Frobenius norm errors for the testing dataset. The average of each distribution is denoted by the corresponding colored circle. The CNN based methods all improve over the simple Born approximation.
@@ 5.2 Noise robustness
Noise, arising from practical measurement apparatus errors or environmental factors, often corrupts scattering data, posing significant challenges to accurately reconstructing scatterers. We trained the CNNs using noise free data (apart from numerical error), so it is important to determine how the various CNN models cope with out-of-distribution data corresponding to far-field patterns corrupted by noise. In this context, following [9], we consider a far-field matrix affected by noise, with entries given by
@@
where $\delta$ controls the strength of the noise and $\mathcal{E}_{i,j}$ is the (complex-valued) noise for incoming wave direction $\varphi_{i}$ and measurement direction $\theta_{j}$. To model the noise, we assume that $\mathcal{E}_{i,j}$ is sampled from a complex Gaussian random variable with zero mean and unit variance:
$\mathcal{E}_{i,j}\sim\mathcal{C}\mathcal{N}(0,1)\iff\Re(Z)\perp\!\!\!\perp\Im(Z)\text{ and }\Re(Z)\sim\mathcal{N}(0,1/2),\>\Im(Z)\sim\mathcal{N}(0,1/2).$
[…]
@@ 5.3 Absorption
Our CNNs are trained on data from dissipation free media ($\Im(\eta)=0$). In many practical applications, the media exhibits varying degrees of dissipation or loss [11], characterized by $\Im(\eta(\textbf{x}))>0$. Since we did not train using far field data from dissipative scatterers, it is important to determine if we can predict the real part of $\eta$ even when the scattering media are slightly dissipative. To investigate such scenarios, we consider a case in which $\eta$ is perturbed by an unknown, small absorption term. Specifically, we use the following imaginary $\eta$ to generate far field test data:
$\eta(\textbf{x})=\eta_{1}(\textbf{x})+\mathrm{i}\mkern 1.0mu\alpha\eta_{2}(\textbf{x}),$
@@
The objective of this experiment is to evaluate the models’ ability to accurately reconstruct $\Re(\eta(\textbf{x}))$ despite the presence of the absorption term in the true scatterer. In particular, we observe that the CNNB architecture can effectively handle the exponential decay in the scattered wave caused by absorption, making it a promising candidate for practical applications where absorption effects, though small, are present. In contrast, the BCNN and CNN networks fail to preserve the shape of the scatterer.
@@ 5.4 Increased scatterer complexity
We next evaluate the generalization capabilities of the CNN models when applied to more complex out-of-distribution scatterers. In particular, we have reconstructed the following:
@@
    We first consider a high-contrast U-shaped resonant structure as shown in Figure  9. The inverse Born scheme has artifacts and does not clearly show the U. While all three CNN models greatly improve the Born approximation, the CNNB model preserves the shape of the U with minimal background artifacts. Meanwhile, the BCNN model constructs the scatterer with distortion in the surrounding media and the CNN model is blurred.
@@
    Our next example tests if it is possible to observe the inner structure of an object. We consider a high-contrast annulus with an elliptic inclusion as shown in Figure  10. The direct use of the inverse Born scheme does not reveal the inner ellipse. All three CNN models perform relatively similarly in obtaining the shape, but the CNNB model acquires the closest approximation to the true shape of the inner ellipse and outer annulus. The BCNN model well approximates the true value of $\eta$ in the annulus but suffers from distortion in the background media.
@@
    In biomedical applications, it is desirable to determine the refractive index of structures within other structures. In this example, we construct an scatterer akin to the Shepp-Logan phantom (restrictions on our mesh generator for the forward problem prevented us using the Shepp-Logan phantom itself). High-contrast scatterers are placed within a high-contrast ring as shown in Figure  11. With appropriate regularization, the inverse Born scheme correctly images (though under-approximates) the outer boundary of the scatterer, but fails to image structures inside. All CNN based methods improve the Born approximation, though the BCNN model well-approximates both the outer ring and the separation of the three internal scatterers.
@@
    Sharp corners can often induce resonance-like behavior in scattering. We consider a shape possessing both sharp corners and an interior of air invisible to the Born approximation. All three CNN models perform comparatively similarly, but the CNNB and BCNN models better approximate the shape of the internal region in comparison to the CNN and Born models.
  5. 5.
    We finally consider a case in which two circular scatterers overlap and are separated by a higher contrast in the overlap. All three CNN models perform comparatively the same, though the BCNN model preserves the shape of the circles and accurately approximates the interior region the best.
@@
Figure 10: The reconstruction of a high-contrast annulur scatterer. All three CNN based schemes are markedly better than inverse Born alone. With the CNN approach, the interior of the annulus can be reconstructed.
Figure 11: The reconstruction of a Shepp-Logan-like phantom. Here the scattering ellipses are surrounded by a thin high contrast annulus of varying thickness. The Born approximation can detect this ring, but fails to reconstruct inside. All three CNN based schemes detect the three internal scatterers.
@@
Figure 13: The reconstruction of two overlapping circles (note the change in the colormap scale). In this case the Born approximation correctly reconstructs the low contrast circles but underestimates refractive index of the inner lense shaped structure. Again all three CNN based schemes improve the estimate at the cost of some fidelity on the lower contrast regions.
@@ 6 Conclusion
Our examples demonstrate that combining CNNs with the Born approximation can extend the applicability of the method beyond the weak scattering limit. We elected to train using simple shapes (disks), but the trained CNN successfully generalizes to more exotic cases. Because of it’s stability to noise and optimality for in-distribution tests, we prefer the CNNB approach.
Much more work needs to be done to extend this demonstration to a realistic biomedical problem like ultrasound tomography. For example, the current implementation of the regularized inverse Born approximate uses simple dense linear algebra to evaluate the necessary inverse. Fast methods should be used, such as the specialized neural network of [25] or the low-rank method of [24]. In addition, for specific applications, other simple training data could be considered (for example if the goal is to image a network).
A particularly interesting direction for further work is to apply the method to more exotic measurement scenarios. For example, the case where both transmitters and receivers are on one side of the object, or when their is missing data from certain sectors.

### v2

@@ 5.1 In-distribution performance
Following the discussion in Section 3.1, we generated 4,000 in-distribution scatterers to evaluate model performance. Table 2 summarizes the average $\ell^{2}$ and $\ell^{1}$ errors of the test scatterer reconstructions, and Figure 4 shows the actual profile of one example of a test scatterer together with the reconstructions.
To make clear that the CNN based models outperform the inverse Born approximation by itself, Figure 5 visualizes the error distribution for the test samples. The results show that the three models–CNN, CNNB, and BCNN–significantly outperform the Born approximation in terms of accuracy, reducing the relative $\ell^{2}$ error by almost 50%.
@@
Table 2: Average relative $\ell^{2}$ and $\ell^{1}$ error in the reconstruction of the contrast $\mu$ expressed as a percentage on the test dataset of $4,000$ scatterers. The rows labeled Born show the results of using the regularized inverse Born approximation. The CNNB model performs optimally in both $\ell^{2}$ and $\ell^{1}$.
$\ell^{2}$ error (%) $\ell^{1}$ error (%)
Born ($\gamma=1$) 67.0378 111.8617
Born ($\gamma=0.1)$ 74.5836 152.5471
CNN 35.0081 76.5449
CNNB 34.2960 69.4069
BCNN 36.1794 91.2601
Figure 4: An example of the reconstructions of three circular scatterers determined by each model. Here we show density plots of the exact and reconstructed values of $\eta$. The ground truth is shown in the upper left panel. All CNN based methods are much more accurate than the inverse Born approximation since, in this example, the scattering is strong. CNNB (bottom left) shows the least speckle in the background.
Figure 5: The distribution of relative $\ell^{2}$ (left) and $\ell^{1}$ (right) norm errors for the test dataset. The average of each distribution is denoted by the corresponding colored circle. The CNN based methods all improve over the simple Born approximation.
@@ 5.2 Noise robustness
Noise, arising from practical measurement apparatus errors or environmental factors, often corrupts scattering data, posing significant challenges to accurately reconstructing scatterers. We trained the CNNs using noise free data (apart from numerical error), so it is important to determine how the various CNN models cope with out-of-distribution data corresponding to far field patterns corrupted by noise. In this context, following [10], we consider a far field matrix affected by noise, with entries given by
@@
where $\delta$ controls the strength of the noise and $\mathcal{E}_{i,j}$ is the (complex-valued) noise for incoming wave direction $\varphi_{i}$ and measurement direction $\theta_{j}$. To model the noise, we assume that $\mathcal{E}_{i,j}$ is sampled from a univariate complex standard normal distribution ${\cal CN}(0,1)$ [1, Def 2.1].
@@ 5.3 Absorption
Our CNNs are trained on data from dissipation free media ($\Im(\eta)=0$). In many practical applications, the media exhibits varying degrees of dissipation or loss [12], characterized by $\Im(\eta)>0$. Since we did not train using far field data from dissipative scatterers, it is important to determine if we can predict the real part of $\eta$ even when the scattering media are slightly dissipative. The objective of this experiment is to evaluate the CNN models’ ability to accurately reconstruct $\Re(\eta(\textbf{x}))$ despite the presence of absorption in the true scatterer.
To investigate such scenarios, we consider the case in which $\eta$ inside the circle is perturbed by an unknown, small absorption term. Specifically, we use the following complex-valued $\eta$ to generate far field test data:
$\eta^{\alpha}\mathrel{\mathop{:}}=\eta_{0}+\mathrm{i}\mkern 1.0mu\alpha$ (23)
@@
The results are shown in Figure 7. We observe that the CNNB architecture can effectively handle the exponential decay in the total field caused by absorption in the scatterer, making it a promising candidate for practical applications where absorption effects, though small, are present. In contrast, the BCNN and CNN networks are less stable to absorption and fail to preserve the shape of the scatterer for $\alpha=10$.
@@ 5.4 Increased scatterer complexity
We next evaluate the generalization capabilities of the CNN models when applied to more complex out-of-distribution scatterers. The scatterers are as follows:
@@
    A U-shaped scatterer chosen to resonate with the incident field (see top left panel in Figure 8). As we shall see, for this shape, the opening of the U is difficult to image using the inverse Born method.
@@
    A high-contrast annulus or ring inside a circle with an elliptic inclusion as shown in Figure 9. The goal with this example is to test if the CNN based schemes can detect the inner wall, even though they were trained on scattering by simple disks.
@@
    In potential biomedical applications, it is desirable to determine the refractive index of structures within other structures. In this example, we construct a scatterer akin to the Shepp-Logan phantom. High-contrast scatterers are placed within a high-contrast ring as shown in the top left hand panel of Figure  10. Restrictions on our mesh generator for the forward problem prevented us using the Shepp-Logan phantom itself. We also note that the values of $\eta$ used here are not the same as for the real Shepp-Logan phantom.
@@
    The CNN based methods are trained on smooth scatterers. Our last test uses a scatterer made of rectangles and an inner low contrast region (see top left panel in Figure 11).
The results of running the various reconstruction algorithms are summarized in Table 3 and in Figures 811 and are discussed next:
Table 3: Relative $\ell^{2}$ error in the reconstruction of $\mu$ as percentages for the inverse Born approximation and the three CNN models for the complex scatterers in Section 5.4. The best reconstruction for each problem is in boldface.
Relative $\ell^{2}$ Error (%)
Method U Ring Shepp Rectangles
Born ($\gamma=0.1$) 92.72 106.26 104.44 87.68
Born ($\gamma=1$) 88.95 87.49 92.76 81.60
CNN 67.63 44.41 63.42 43.66
CNNB 60.53 41.23 60.03 39.28
BCNN 66.28 39.76 60.57 43.03
  1. 1.
    The results of reconstructing the high-contrast U-shaped resonant structure are shown in Figure  8 and Table 3. The inverse Born scheme exhibits artifacts and does not clearly show the U. While all three CNN models greatly improve on the Born approximation, the CNNB model preserves the shape of the U with minimal background artifacts. Meanwhile, the BCNN model constructs the scatterer with distortion in the surrounding media and the CNN model is blurred. The relative error in $\ell^{1}$ or $\ell^{2}$ conform that CNNB is best in this case.
  2. 2.
    Our second example, the high contrast ring scatterer, tests if it is possible to observe the inner structure of an object. As shown in Figure 9, the direct use of the inverse Born scheme does not reveal the inner ellipse. All three CNN models improve of the inverse Born, and perform relatively similarly in obtaining the shape. In this case, the BCNN model best approximates the true value of $\eta$ in the annulus as shown in Table 3.
  3. 3.
    Concerning the modified Shepp-Logan phantom, we see from Figure 10 and Table 3 that with appropriate regularization, the inverse Born scheme correctly images (though under-approximates) $\eta$ in the outer boundary of the scatterer, but fails to image structures inside. All CNN based methods improve over the inverse Born approximation, though the BCNN model well-approximates both the outer ring and the separation of the three internal scatterers, CNNB gives the best quantitative reconstruction.
  4. 4.
    We finally consider the shape constructed from several rectangles, possessing both sharp corners and an interior of air as shown in Figure 11 and Table 3. We see that the inner inclusion is invisible to the Born approximation, and the high contrast block is not well approximated. All three CNN models perform comparatively similarly, but the CNNB and BCNN models better approximate the shape of the internal region in comparison to the CNN and Born models. Quantitatively CNNB performs best. The presence of corners does not badly impact the CNN based models.
@@
Figure 9: The reconstruction of a high-contrast annular scatterer. All three CNN based schemes are markedly better than inverse Born alone. With the CNN approach, the interior of the annulus can be reconstructed.
Figure 10: The reconstruction of a simplified Shepp-Logan-like phantom. Here the scattering ellipses are surrounded by a thin high contrast annulus of varying thickness. The Born approximation can detect this ring, but fails to reconstruct inside. All three CNN based schemes detect the three internal scatterers.
@@ 6 Conclusion
Our examples demonstrate that combining CNNs with the Born approximation can extend the applicability of the method beyond the weak scattering limit. All three CNN models can reveal the inner structure of high contrast objects better than the Born approximation. Moreover, a CNN combined with the inverse Born model always improves scatterer reconstruction in comparison to a pure CNN model. Because of it’s stability to noise and optimality for almost every test, we prefer the CNNB approach.
It is notable that we elected to train using simple shapes (disks), but the trained CNN models successfully generalize to more exotic cases.
Much more work needs to be done to extend this demonstration to a realistic biomedical problem like ultrasound tomography. For example, the current implementation of the regularized inverse Born approximate uses simple dense linear algebra to evaluate the necessary inverse. Fast methods should be used, such as the specialized neural network of [26] or the low-rank method of [25]. In addition, for specific applications, other simple training data could be considered (for example, if the goal is to image a network of blood vessels).
A particularly interesting direction for further work is to apply the method to more exotic measurement scenarios. For example, the case where both transmitters and receivers are on one side of the object, or when there is missing data from certain angular sectors. Moreover, additional work could include extending the applicability of the model to cases in which the wave number $k$ varies (i.e. multi-frequency data) or when the scattering medium is defined by a function that is not piecewise-constant.